OpenStudy (anonymous):

Solve by completing the square: x^2 - 2x - 24 = 0 A. -4, -6 B. -4, 6 C. 4, -6 D. 4, 6

ganeshie8 (ganeshie8):

start by adding 24 both sides

ganeshie8 (ganeshie8):

x^2 - 2x - 24 = 0 +24 +24 x^2-2x = 24

ganeshie8 (ganeshie8):

x^2-2(1)x = 24 add 1^2 both sides x^2-2(1)x + 1^2 = 24 + 1^2

ganeshie8 (ganeshie8):

Now stare at the left hand side, can does it look like a perfect square ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

good :) write it as a square :- \(a^2-2ab + b^1 = (a-b)^2\)

ganeshie8 (ganeshie8):

x^2-2(1)x + 1^2 = 24 + 1^2 (x-1)^2 = 25 take sqrt both sides

ganeshie8 (ganeshie8):

wat do you get ?

OpenStudy (anonymous):

I don't know where to go from there.

ganeshie8 (ganeshie8):

its okay :) its easy actually, once u see it done, u wil get it

ganeshie8 (ganeshie8):

\(\large (x-1)^2 = 25\) take sqrt both sides \(\large \sqrt{(x-1)^2} = \sqrt{25}\)

ganeshie8 (ganeshie8):

square, and sqrt cancel out on left side

ganeshie8 (ganeshie8):

\(\large \sqrt{(x-1)^2} = \sqrt{25}\) \(\large x-1 = \pm 5\) that gives, \(\large x-1 = + 5 \) (or) \(\large x-1 = -5\) \(\large x = 6 \) (or) \(\large x = -4\)

ganeshie8 (ganeshie8):

check if my calculations were correct...

OpenStudy (anonymous):

Okay. I will look through this. Thank you.

ganeshie8 (ganeshie8):

np :)